# MIDAS: Multi-level Intent, Domain, And Slot Knowledge Distillation for Multi-turn NLU

### v2

@@ G.1 Quantitative analysis
GPT4o with demonstration 0.7614 0.7510 0.8525 0.7132 0.7941 0.7051
@@
Table 10: The comparison of the proposed models with prompt tuning methods using Large Language Models. ID, SF and DC indicate intent detection, slot filling and domain classification, respectively, as mentioned in Section 4.2. ACC and F1 stand for accuracy and micro F1, respectively, and scores in bold indicate leadership among the metrics.
We measured the performance using the zero-shot prompt method to compare performance with Large Language Model (LLM). The LLM LLaMa, Gemma, and GPT3.5 were utilized. The prompt for Intent Detection (ID) and Domain Classification (DC) was given as follows: In this task, you are given a dialogue. Your job is to classify the following dialogue into one of the {number of classes} different {intents or domains}. The {intents or domains} are: {name of classes}. Input: [{input}]. Output(only output the {intent or domain}):. The prompt for Slot Filling (SF) is: In this task, you are given a dialogue. Your job is to classify the following dialogue into one of the {number of classes} different slots. The slots are: {name of classes}. Input: [{input}]. Output(Only output slot types. And the slot types should be output as a list without any explanation):.
Table 10 shows the experimental results of each baseline with the performance of our best model. We can see that GPT3.5 performs best on all tasks in the zero-shot testing, but still falls significantly short of our model’s performance. In the ID and SF tasks, the performance of LLaMa is significantly worse than that of Gemma and GPT. This suggests that factors such as architecture, training data, and training methods also impact LLM performance, in addition to the number of parameters.
@@
G.2 Qualitative analysis
Figure 2: Two examples for qualitative analysis: (a) shows the results on the M2M dataset, and (b) shows the results on the MWOZ dataset. Each example shows the results when MIDAS matches the ground truth. The three cells below each example display the type lists for slot, intent, and domain and red text in each column of the results table indicates errors.
@@ G.2 Qualitative analysis
In the qualitative analysis, we focus on two representative LLMs, Gemma-7b and GPT3.5, as shown in Figure 2. From the M2M conversation as shown in Figure 2-(a), we found that both LLMs can not predict slot types based on the context. For example, GPT3.5 predicts “Michelin/B-RN rated/I-Rating” instead of “Michelin/B-Rating rated/I-Rating”. Except for the wrong understanding of the conversation, we found that both LLMs can not follow the prompt all the time. For example, both LLMs do not predict the slot type for each token, where the missing predictions are represented by “N”. From the Multi-Domain Wizard-of-Oz 2.2 (MWOZ) conversation as shown in Figure 2-(b), we can see that both LLMs can not make predictions in terms of the whole conversation, resulting the conflicts of the predictions of the domains and intents. For example, GPT3.5 predicts “Book Train” after “Book Restaurant” and Gemma-7b predicts “Find Restaurant” after “Book Restaurant”. Another example is that both LLMs failed to predict the domain “Restaurant” of the last turn dialogue, even the Gemma-7b already predicted the intent as “Find Restaurant”.
Appendix H Implementation details
H.1 Experiment hyperparameters
Hyper-parameter   Value in Fine-tuning   Value in Knowledge Distillation
Learning Rate 5e-5 5e-5
Batch Size 32 32
Warm-up Steps 10% of Max epoch 10% of Max epoch
Mex epoch 3 100
Stop Strategy Max Epoch Early Stopping on validation loss
Stop Patience - 10
Optimizer AdamW AdamW
Optimizer Weight Decay 1e-2 1e-2
Optimizer Betas 0.9, 0.999 0.9, 0.999
Margin in $\mathcal{L}_{rel}$ 0.2
Norm in $\mathcal{L}_{rel}$ 2
$\mathcal{F}_{D}$ in $\mathcal{L}_{rel}$ L2-Norm
Similarity in $\mathcal{L}_{sim}$ Cosine Similarity
Max Token Length 512 512
Table 11: The hyper-parameters used in our experiments.
     Model    Task   Learning Rate   Accuracy
M2M
LLaMa2-7b ID 0.0005 0.9121
0.00005 0.9392
0.000005 0.9093
SF 0.0005 0.9696
0.00005 0.9833
0.000005 0.9349
DC 0.0005 0.8895
0.00005 0.8804
0.000005 0.8375
Gemma-7b ID 0.0005 0.9204
0.00005 0.9357
0.000005 0.9102
SF 0.0005 0.9693
0.00005 0.9816
0.000005 0.9429
DC 0.0005 0.8799
0.00005 0.8840
0.000005 0.7890
MWOZ
LLaMa2-7b ID 0.0005 0.8021
0.00005 0.8403
0.000005 0.7952
SF 0.0005 0.9776
0.00005 0.9912
0.000005 0.9740
DC 0.0005 0.8411
0.00005 0.8702
0.000005 0.7026
H.2 Model details
  BERT   RoBERTa   LLaMa   Student
Architecture Encoder Encoder Decoder Encoder
Parameters 110M 125M 7B 58M
Layers 12 12 32 6
Heads 12 12 32 8
Hidden Dim. 768 768 4096 768
Feed Forward Dim. 3072 3072 11008 2048
Dropout Rate 0.1 0.1 0.0 0.3
Rank of LoRA - - 64
Alpha of LoRA - - 16
Dropout of LoRA - - 0.1
Table 13: The details of the models used in our work.
H.3 Hardware information

### v3

@@ D.2.1 Quantitative analysis
GPT4$\dagger$ 0.7614 0.7510 0.8525 0.7132 0.7941 0.7051
QWen2-7B-Ins.$\dagger$ 0.5459 0.278 0.1532 0.1192 0.6416 0.6541
LLaMa3.1-8b-Ins.$\dagger$ 0.6422 0.2715 0.6276 0.5412 0.5973 0.5076
@@
Table 14: The comparison of the proposed models with prompt tuning methods using Large Language Models. $ID$, $SF$ and $DC$ indicate intent detection, slot filling and domain classification, respectively, as mentioned in Section 4.3. ACC and F1 stand for accuracy and micro F1, respectively, and scores in bold indicate leadership among the metrics. $\dagger$ refers to the application of 3-shot learning in the prompt with LLMs.
We measured the performance using the zero-shot prompt method to compare performance with Large Language Model (LLM). The LLM LLaMa2Touvron et al. (2023), LLaMa3.1Dubey et al. (2024), GemmaTeam et al. (2024), QWen2Bai et al. (2023), GPT3.5 and GPT4oBrown et al. (2020), were utilized. Notably, we also tested 3-shot learning on GPT4o, QWen2 and LLaMa3.1. The prompt for each task are shown in Table 15.
Table 14 presents the experimental results of each baseline, compared with the performance of our best model. Notably, GPT4o 3-shot learning achieves the best results in all tasks, except for the DC task on the M2M dataset, though it still falls significantly short of our model’s performance. In the ID and SF tasks, LLaMa’s performance is markedly lower than that of Gemma and GPT, highlighting that factors such as architecture, training data, and training methods, beyond just the number of parameters, also influence LLM performance.
@@
Task Prompt
ID Definition: In this task, you are given a dialogue. Your job is to classify the following dialogue into one of the fifteen different intents. The intents and examples are:
Intent: greeting Dialogues: 1. hi , i want to make a restaurant reservation . 2. hi , i want to make a reservation for 6 pm . 3. hi ! can you book me a restaurant reservation ? Intent: request Dialogues: 1. okay , where do you want to go , and how many people will there be? …
Input: [{input}]. Output(only output the intent):
SF In the task of slot filling, the B-, I-, and O- prefixes are commonly used to annotate slot types, indicating the boundaries and types of slots. These labels typically represent:
B- (Begin): Signifies the beginning of a slot, marking the start of a new slot. I- (Inside): Represents the interior of a slot, indicating a continuation of the slot. O (Outside): Denotes parts of the input that are not part of any slot.
For instance, in a sentence where we want to label a "date" slot, words containing date information might be tagged as "B-date" (indicating the beginning of a date slot), followed by consecutive words carrying date information tagged as "I-date" (indicating the continuation of the date slot), while words not containing date information would be tagged as "O" (indicating they are outside any slot). Here are some examples:
Dialogue: "the sushi boat for 6 .", slot types: [’O’, ’B-restaurant_name’, ’I-restaurant_name’, ’O’, ’B-num_people’, ’O’] …
Definition: In this task, you are given a dialogue. Your job is to classify the words in the following dialogue into one of the twenty-one different slots. The slots are: "B-category", "B-date", "B-location", …, "O". Input: [{input}]. Output(Only output slot types. And the slot types should be output as a list without any explanation):
DC In this task, you are given a dialogue. Your job is to classify the following dialogue into one of the two different intents. The domains and examples are:
Domain: restaurant Dialogues: 1. hi , i want to make a restaurant reservation. 2. a reservation for cheese cake factory for 3 people on next monday. 3. ok , please choose between amarin and sakoon restaurants.
Domain: movie Dialogues: 1. i would like to buy movie tickets for 6:00 pm 2. which movie , and how many tickets do you need ? 3. i need 3 tickets for the movie called a man called love
Input: [{input}]. Output(only output the domain):
Table 15: The prompt we used for each dataset in our experiments.
D.2.2 Qualitative analysis
@@
In the qualitative analysis, we first focus on two representative LLMs, Gemma-7b and GPT3.5, as shown in Figure 3. From the M2M conversation shown in Figure 3-(a), we found that both LLMs can not predict slot types based on context. For example, GPT3.5 predicts “Michelin/B-RN rated/I-Rating” instead of “Michelin/B-Rating rated/I-Rating”. Except for the wrong understanding of the conversation, we found that both LLMs can not follow the prompt all the time. For example, both LLMs do not predict the slot type for each token, where the missing predictions are represented by “N”. From the Multi-Domain Wizard-of-Oz 2.2 (MWOZ) conversation as shown in Figure 3-(b), we can see that both LLMs can not make predictions in terms of the whole conversation, resulting the conflicts of the predictions of the domains and intents. For example, GPT3.5 predicts “Book Train” after “Book Restaurant” and Gemma-7b predicts “Find Restaurant” after “Book Restaurant”. Another example is that both LLMs failed to predict the domain “Restaurant” of the last turn dialogue, even the Gemma-7b already predicted the intent as “Find Restaurant”.
We further analyse the outputs from cutting-edge large language models, including QWen2, LLaMa3.1, and GPT-4o. As outlined in Section 5.7, these models also faced challenges in making use of multi-level knowledge consistently and strictly adhering to prompt instructions. Those examples from both M2M dataset and MultiWOZ dataset are presented from Table 16 to Table 21.
D.3 Complete procedure cases
In this section, we provide a complete question-answer demonstration for each task and dataset, offering a clearer explanation of how we utilize LLMs for testing, as shown from Figure 4 to Figure 9.
Figure 4: Prompt and output for a sample dialogue in MultiWOZ dataset, where the correct prediction is highlighted in green and wrong predictions are highlighted in red. Demonstration means few-shot (3-shot) learning. Compared to GPT4o, our model can correctly classify the intent of the given dialogue as Book Restaurant.
Figure 5: Prompt and output for a sample dialogue in M2M dataset, where the correct prediction is highlighted in green and wrong predictions are highlighted in red. Demonstration means few-shot (3-shot) learning. Compared to GPT4o, our model can correctly classify the intent of the given dialogue as Select.
Figure 6: Prompt and output for a sample dialogue in MultiWOZ dataset, where the correct prediction is highlighted in green and wrong predictions are highlighted in red. Demonstration means few-shot (3-shot) learning. Compared to GPT4o, our model can correctly classify the slot types of the given dialogue.
Figure 7: Prompt and output for a sample dialogue in M2M dataset, where the correct prediction is highlighted in green and wrong predictions are highlighted in red. Demonstration means few-shot (3-shot) learning. Compared to GPT4o, our model can correctly classify the slot types of the given dialogue.
Figure 8: Prompt and output for a sample dialogue in M2M dataset, where the correct prediction is highlighted in green and wrong predictions are highlighted in red. Demonstration means few-shot (3-shot) learning. Both models can correctly classify the domain of the given dialogue as Restrurant.
Figure 9: Prompt and output for a sample dialogue in MultiWOZ dataset, where the correct prediction is highlighted in green and wrong predictions are highlighted in red. Demonstration means few-shot (3-shot) learning. Compared to GPT4o, our model can correctly classify the domain of the given dialogue as Restaurant.
Model    Turn                 Tokens (Slot)     Intent    Domain
Utterance 1 what, are, you, in, the, mood, for, ? - -
2 it, doesn, , t, matter, . - -
3 how, about, the, hudson, room, or, los, altos, grill, ? - -
4 sounds, good, . - -
GT 1 O, O, O, O, O, O, O, O request restaurant
2 O, O, O, O, O, O inform restaurant
3 O, O, B-restaurant_name, I-restaurant_name, I-restaurant_name, O, B-restaurant_name, I-restaurant_name, I-restaurant_name, O select restaurant
4 O, O, O affirm restaurant
MIDAS (Ours) 1 O, O, O, O, O, O, O, O request restaurant
2 O, O, O, O, O, O inform restaurant
3 O, O, B-restaurant_name, I-restaurant_name, I-restaurant_name, O, B-restaurant_name, I-restaurant_name, I-restaurant_name, O select restaurant
4 O, O, O affirm restaurant
GPT4o 1 O, O, O, O, O, O, NaN, NaN request restaurant
2 O, O, O, O, O, NaN negate (O)
3 O, O, O, B-restaurant_name, I-restaurant_name, O, B-restaurant_name, I-restaurant_name, O, NaN select restaurant
4 O, O, O affirm movie
QWen2 1 B-category, NaN, NaN, NaN, NaN, NaN, NaN, NaN request_alts restaurant
2 O, O, NaN, NaN, NaN, NaN negate restaurant
3 B-restaurant_name, I-restaurant_name, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN request_alts restaurant
4 O, O, NaN inform movie
LLaMa3.1 1 O, O, O, O, O, NaN, NaN, NaN request movie
2 O, O, O, O, O, O other movie
3 O, O, O, O, O, O, O, O, O, O affirm movie
4 O, O, O other restaurant
Table 16: A four turns conversation from M2M dataset. GPT4o exhibited issues with generating out-of-domain outputs ‘(O)’ in Turn 2, while all LLMs showed problems in the SF task by failing to follow instructions, resulting in NaN outputs.
Model    Turn                 Tokens (Slot)     Intent    Domain
Utterance 1 what, area, is, your, preference, ? - -
2 orlando - -
3 i, have, the, hudson, room, ,, boats, or, the, nest, ,, which, sounds, better, ? - -
4 agree, on, the, hudson, room - -
GT 1 O, O, O, O, O, O request restaurant
2 B-location inform restaurant
3 O, O, B-restaurant_name, I-restaurant_name, I-restaurant_name, O, B-restaurant_name, O, B-restaurant_name, I-restaurant_name, O, O, O, O, O select restaurant
4 O, O, B-restaurant_name, I-restaurant_name, I-restaurant_name affirm restaurant
MIDAS (Ours) 1 O, O, O, O, O, O request restaurant
2 B-location inform restaurant
3 O, O, B-restaurant_name, I-restaurant_name, I-restaurant_name, O, B-restaurant_name, O, B-restaurant_name, I-restaurant_name, O, O, O, O, O select restaurant
4 O, O, B-restaurant_name, I-restaurant_name, I-restaurant_name affirm restaurant
GPT4o 1 O, O, O, O, O, NaN request restaurant
2 B-location (O) (O)
3 O, O, O, B-restaurant_name, I-restaurant_name, O, O, O, O, O, B-restaurant_name, O, O, O, O select restaurant
4 O, O, B-restaurant_name, I-restaurant_name, NaN inform restaurant
QWen2 1 B-location, NaN, NaN, NaN, NaN, NaN request_alts movie
2 B-location inform movie
3 B-location, B-location, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN request_alts movie
4 B-restaurant_name, I-location, NaN, NaN, NaN select restaurant
LLaMa3.1 1 O, O, O, O, O, NaN greeting movie
2 O confirm movie
3 O, O, O, B-location, O, B-movie, O, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN request_alts movie
4 O, O, O, B-location, I-location thank_you movie
Table 17: A four turns conversation from M2M dataset. GPT4o exhibited issues with generating out-of-domain outputs ‘(O)’ in Turn 2, while all LLMs showed problems in the SF task by failing to follow instructions, resulting in NaN outputs.
Model    Turn                 Tokens (Slot)     Intent    Domain
Utterance 1 how, many, tickets, would, you, like, to, buy, ? - -
2 1 - -
3 what, date, do, you, want, to, go, ? - -
GT 1 O, O, O, O, O, O, O, O, O request movie
2 B-num_tickets inform movie
3 O, O, O, O, O, O, O, O request movie
MIDAS (Ours) 1 O, O, O, O, O, O, O, O, O request movie
2 B-num_tickets inform movie
3 O, O, O, O, O, O, O, O request movie
GPT4o 1 O, O, O, O, O, O, O, NaN, NaN request movie
2 (O) (O) restaurant
3 O, B-date, O, O, O, O, NaN, NaN request restaurant
QWen2 1 B-num_tickets, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN request_alts movie
2 B-category inform movie
3 B-date, NaN, NaN, NaN, NaN, NaN, NaN, NaN request_alts restaurant
LLaMa3.1 1 O, O, O, O, O, O, B-num_tickets, O, NaN request restaurant
2 B-category request movie
3 O, O, O, O, O, O, O, O request movie
Table 18: A three turns conversation from M2M dataset. GPT4o exhibited issues with generating out-of-domain outputs ‘(O)’ in Turn 2, while all LLMs showed problems in the SF task by failing to follow instructions, resulting in NaN outputs. LLaMa3.1 failed to unify multi-level knowledge, resulting in predicting the domain as ‘restaurant’ even after identifying ‘B-num_tickets’
Model    Turn                 Tokens (Slot)     Intent    Domain
Utterance 1 1, pm, is, unavailable, again, ,, please, choose, from, 6, pm, or, 6.30, pm, . - -
2 no, ,, thank, you - -
3 please, provide, the, theatre, name, ,, date, ,, and, number, of, tickets - -
4 the, theatre, name, is, aquarius, and, i, need, 1, ticket, for, friday, . - -
5 what, movie, would, you, like, to, see, ? - -
GT 1 B-time, I-time, O, O, O, O, O, O, O, B-time, I-time, O, B-time, I-time, O negate movie
2 O, O, O, O negate movie
3 O, O, O, O, O, O, O, O, O, O, O, O request movie
4 O, O, O, O, B-theatre_name, O, O, O, B-num_tickets, O, O, B-date, O inform movie
5 O, O, O, O, O, O, O, O request movie
MIDAS (Ours) 1 B-time, I-time, O, O, O, O, O, O, O, B-time, I-time, O, B-time, I-time, O negate movie
2 O, O, O, O negate movie
3 O, O, O, O, O, O, O, O, O, O, O, O request movie
4 O, O, O, O, B-theatre_name, O, O, O, B-num_tickets, O, O, B-date, O inform movie
5 O, O, O, O, O, O, O, O request movie
GPT4o 1 B-time, I-time, O, O, O, O, O, O, O, B-time, I-time, O, O, B-time, I-time negate movie
2 O, O, O, NaN negate (O)
3 O, O, O, B-theatre_name, I-theatre_name, O, B-date, O, O, B-num_tickets, O, NaN request movie
4 O, B-theatre_name, I-theatre_name, O, O, O, O, B-num_tickets, O, O, B-date, O, NaN inform movie
5 O, B-movie, O, O, O, O, O, NaN request movie
QWen2 1 B-time, I-time, I-time, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN request_alts (O)
2 O, O, NaN, NaN negate restaurant
3 B-theatre_name, B-date, B-num_tickets, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN request_alts movie
4 B-theatre_name, B-date, B-num_tickets, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN request movie
5 B-movie, NaN, NaN, NaN, NaN, NaN, NaN, NaN request_alts movie
LLaMa3.1 1 O, O, O, B-time, I-time, O, O, O, O, O, O, O, O, O, O request restaurant
2 O, O, NaN, NaN thank_you restaurant
3 B-theatre_name, I-date, B-num_tickets, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN other movie
4 B-theatre_name, I-theatre_name, O, O, O, B-num_tickets, (O)I-num_tickets, O, B-date, I-date, O, NaN, NaN request movie
5 O, O, O, O, O, O, O, O request movie
Table 19: A five turns conversation from M2M dataset. GPT4o, QWen2 and LLaMa3.1 exhibited issues with generating out-of-domain outputs ‘(O)’ in Turn 2, Turn 1 and Turn 4 respectively, while all LLMs showed problems in the SF task by failing to follow instructions, resulting in NaN outputs.
Model    Turn                 Tokens (Slot)     Intent    Domain
Utterance 1 i, choose, the, ashley, hotel., what, is, their, address,, please? - -
2 no,, i, just, need, the, address. - -
3 hey, i, am, looking, for, a, train, from, cambridge, to, bishops, stortford., mind, helping, out? - -
4 i, want, to, leave, on, monday, and, arrive, by, 18:45. - -
5 could, you, give, me, the, travel, time, and, price, of, that, train, please? - -
6 i, also, want, a, cheap, chinese, restaurant. - -
7 id, like, to, be, in, the, centre, please. - -
8 yes., monday,, 8, people,, 10:30. - -
9 i, am, planning, a, trip, in, cambridge - -
GT 1 O, O, O, B-hotel-name, I-hotel-name, O, O, O, O, O find_hotel hotel
2 O, O, O, O, O, O find_hotel hotel
3 O, O, O, O, O, O, O, O, O, O, O, O, O, O, O find_train train
4 O, O, O, O, O, O, O, O, O, B-train-arriveby find_train train
5 O, O, O, O, O, O, O, O, O, O, O, O, O find_train train
6 O, O, O, O, O, B-restaurant-food, O find_restaurant restaurant
7 O, O, O, O, O, O, O, O find_restaurant restaurant
8 O, O, O, O, B-restaurant-booktime book_restaurant restaurant
9 O, O, O, O, O, O, O find_train train
MIDAS (Ours) 1 O, O, O, B-hotel-name, I-hotel-name, O, O, O, O, O find_hotel hotel
2 O, O, O, O, O, O find_hotel hotel
3 O, O, O, O, O, O, O, O, O, O, O, O, O, O, O find_train train
4 O, O, O, O, O, O, O, O, O, B-train-arriveby find_train train
5 O, O, O, O, O, O, O, O, O, O, O, O, O find_train train
6 O, O, O, O, O, B-restaurant-food, O find_restaurant restaurant
7 O, O, O, O, O, O, O, O find_restaurant restaurant
8 O, O, O, O, B-restaurant-booktime book_restaurant restaurant
9 O, O, O, O, O, O, O find_train train
GPT4o 1 O, O, O, B-hotel-name, I-hotel-name, O, O, O, O, O find_hotel hotel
2 O, O, O, O, O, NaN find_police police
3 O, O, O, O, O, O, O, (O)B-train-departure, O, (O)B-train-destination, (O)I-train-destination, O, O, O, O find_train train
4 O, O, O, O, O, O, B-train-arriveby, I-train-arriveby, NaN, NaN find_train train
5 O, O, O, O, O, O, O, O, O, O, O, O, O find_train train
6 O, O, O, O, B-restaurant-food, I-restaurant-food, O find_restaurant restaurant
7 O, O, O, O, O, O, O, NaN find_hotel hotel
8 O, O, O, B-restaurant-booktime, O book_restaurant train
9 O, O, O, O, O, B-bus-destination, NaN find_attraction attraction
QWen2 1 B-hotel-name, I-hotel-name, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN find_hotel hotel
2 O, O, NaN, NaN, NaN, NaN find_hotel hotel
3 (O)B-train-destination, (O)B-train-origin, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN book_train train
4 B-train-leaveat, (O)I-train-arrivevi, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN book_train hotel
5 O, O, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN find_train taxi
6 I-restaurant-food, NaN, NaN, NaN, NaN, NaN, NaN find_restaurant restaurant
7 I-hotel-stars, NaN, NaN, NaN, NaN, NaN, NaN, NaN find_hotel hotel
8 B-train-leaveat, I-restaurant-booktime, NaN, NaN, NaN book_train hotel
9 (O)B-city, B-hotel-name, NaN, NaN, NaN, NaN, NaN find_attraction hotel
LLaMa3.1 1 O, O, O, O, B-hotel-name, O, O, O, O, O find_hotel hotel
2 B-restaurant-name, I-restaurant-name, O, O, NaN, NaN find_hotel restaurant
3 B-train-leaveat, O, (O)B-train-destination, O, O, O, O, O, O, O, O, O, O, O, O find_train train
4 B-train-leaveat, O, O, B-train-arriveby, O, O, O, O, O, O book_train train
5 B-train-arriveby, B-train-leaveat, O, O, O, O, O, O, O, O, O, O, O find_train train
6 O, O, O, B-restaurant-food, O, O, O find_restaurant restaurant
7 O, O, O, O, O, O, O, O find_taxi attraction
8 B-attraction-name, I-attraction-name, O, O, O find_restaurant bus
9 O, O, O, O, O, O, O find_attraction restaurant
Table 20: A nine turns conversation from MultiWOZ dataset. All LLMs showed problems generating out-of-domain outputs ‘(O)’ and problems in the SF task by failing to follow instructions, resulting in NaN outputs.
Model    Turn                 Tokens (Slot)     Intent    Domain
Utterance 1 yes, i, am, looking, for, a, place, to, stay, in, cambridge, that, is, 3, stars, and, expensive., can, you, help, me? - -
2 do, those, both, have, 3, star, ratings, and, are, expensive? - -
3 do, they, have, free, parking? - -
4 not, today, thanks., im, also, want, to, find, a, cinema, in, the, west, part, of, town. - -
5 are, there, any, colleges, in, the, west, that, i, could, visit, instead? - -
6 what, is, the, one, that, is, free?, can, i, get, the, phone, number, and, postcode? - -
GT 1 O, O, O, O, O, O, O, O, O, O, O, O, O, O, O, O, O, O, O, O, O find_hotel hotel
2 O, O, O, O, O, O, O, O, O, O find_hotel hotel
3 O, O, O, O, O find_hotel hotel
4 O, O, O, O, O, O, O, O, O, O, O, O, O, O, O, O find_attraction attraction
5 O, O, O, O, O, O, O, O, O, O, O, O find_attraction attraction
6 O, O, O, O, O, O, O, O, O, O, O, O, O, O, O find_attraction attraction
MIDAS (Ours) 1 O, O, O, O, O, O, O, O, O, O, O, O, O, O, O, O, O, O, O, O, O find_hotel hotel
2 O, O, O, O, O, O, O, O, O, O find_hotel hotel
3 O, O, O, O, O find_hotel hotel
4 O, O, O, O, O, O, O, O, O, O, O, O, O, O, O, O find_attraction attraction
5 O, O, O, O, O, O, O, O, O, O, O, O find_attraction attraction
6 O, O, O, O, O, O, O, O, O, O, O, O, O, O, O find_attraction attraction
GPT4o 1 O, O, O, O, O, O, O, O, O, O, O, O, O, B-hotel-stars, I-hotel-stars, O, O, O, O, O, NaN find_hotel hotel
2 O, O, O, O, B-hotel-stars, I-hotel-stars, O, O, O, NaN find_restaurant restaurant
3 O, O, O, O, O find_hotel hotel
4 O, O, O, O, O, O, O, O, O, O, O, O, O, O, O, O find_attraction attraction
5 O, O, O, O, O, O, O, O, O, O, NaN, NaN find_attraction attraction
6 O, O, O, O, O, O, O, O, O, O, O, O, O, O, O find_attraction attraction
QWen2 1 B-hotel-name, B-hotel-stars, (O)B-price-range, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN find_hotel hotel
2 B-hotel-stars, B-hotel-type, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN find_hotel restaurant
3 O, O, NaN, NaN, NaN find_hotel restaurant
4 B-attraction-name, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN find_attraction attraction
5 O, O, O, O, O, O, O, O, O, O, O, O find_attraction attraction
6 O, O, O, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN, NaN find_attraction restaurant
LLaMa3.1 1 B-hotel-stars, I-hotel-stars, O, B-hotel-type, O, B-hotel-stars, I-hotel-stars, O, B-hotel-type, O, B-hotel-stars, I-hotel-stars, O, B-hotel-type, O, NaN, NaN, NaN, NaN, NaN, NaN book_hotel hotel
2 O, O, O, O, O, O, O, O, O, O find_hotel restaurant
3 O, O, NaN, NaN, NaN find_hotel hotel
4 O, O, O, O, O, O, O, O, O, O, O, O, O, O, O, O find_attraction restaurant
5 O, O, O, O, O, O, O, O, O, O, O, O find_attraction attraction
6 O, O, O, O, O, O, O, O, O, O, O, O, O, O, O find_hotel hotel
Table 21: A six turns conversation from MultiWOZ dataset. QWen2 exhibited issues with generating out-of-domain outputs ‘(O)’ in Turn 1, while all LLMs showed problems in the SF task by failing to follow instructions, resulting in NaN outputs.